# Amorphous silicon structures generated using a moment tensor potential and the activation relaxation technique nouveau

### v1

@@ III.5 Potential energy
Fig. 11: Excess energy between ART-MTP configurations (a-Si) and crystalline silicon (c-Si) as a function of the number of accepted events. A total of 4 events per atom was required to achieve 100% 4-fold coordination in 216-R2 a-Si model.
@@
Fig. 11 shows the relationship between energy decrease and the number of events for the model 216-R2. As can be seen, the energy decreases as the number of events increase. A detailed analysis of the events and key parameters is provided in the Nature of the Relaxation section. The experimental excess energy interval is indicated by the light red shaded region in the figure for reference. The inset of Fig. 11 shows the excess energy per atom for each of our models. These values are also presented in table 4. All models fall within the experimental excess energy interval 81, as detailed in the inset. Notably, the 216-R2 model, which initially contained over 70% coordination defects (see Fig. 2), was efficiently relaxed to achieve 100% 4-fold coordination and the lowest excess energy of 0.17 eV/atom with just 4 events per atom. To our knowledge, with the exception of a few Keating-potential-based WWW models 10, 78, 23, no other method, including DFT 84, has yielded a model of a-Si with no coordination defects.
@@ III.6 Density
We computed the density of our models at 0 K. The final density was obtained by running a 50 ps molecular dynamics simulation, cooling the system from 300 K to 0.5 K, and then holding it at 0.5 K for an additional 50 ps, while barostatting the system at 0 Pa. In separate simulations, we also equilibrated each of the aforementioned models at 300 K and 1 atm for 100 ps. The results are presented in table 4. The ART-MTP a-Si density values at 0 K are 0.71% to 1.94% larger than the experimental density measured at room temperature. ART-MTP a-Si density at 300 K and 1 atm is characterized by density values 1.18% to 2.43% larger than experimental densities. The MTP potential is trained using a generalized gradient approximation exchange-correlation functional, which is known to overestimate the density of crystalline silicon by 2 % 85. Retraining the MTP on a higher-accuracy exchange-correlation functional, such as the restored regularized strongly constrained and appropriately normed meta-generalized gradient approximation 86 is likely to solve this small discrepancy with experimental values.
@@ III.7 Nature of the relaxation
The primary goal is not only to generate high-quality amorphous structures but also to understand the underlying physics of the relaxation process in amorphous systems over time. Thus, to understand the relaxation process in our amorphous system, we analyze simulation data, focusing on key parameters such as barrier distribution, atomic displacements, active atoms, and energy decrease, considering only accepted events as detailed in Fig. 12. These parameters can help provide insights into the nature and mechanism of the relaxation process. The data were collected under normal deformation, excluding shear deformation during the simulation. We first examine the correlation between atomic displacement and energy barriers. A clear correlation is observed: larger displacements are typically associated with higher energy barriers and more active atoms, as shown in Fig. 12 (a). To further validate this, we calculate the correlation between the square of the atomic displacement and the energy barriers. According to Hooke’s law, we expect a quadratic relationship between energy and displacement. When we plot energy against the square of the displacement, we expect a linear relationship, which our results confirm, yielding a strong correlation. Next, we find no significant correlation between the energy barriers and the energy gain, nor between energy gain and the number of active atoms, which is consistent with previous studies 87. Furthermore, no correlation exists between the number of atoms displaced up to the saddle point and the final state. As illustrated in Fig. 12(d), a correlation exists between energy barriers and the number of active atoms, with higher energy barriers typically involving a greater number of active atoms. On average, the number of active atoms varies from 10 to 30, with a mean of around 20 atoms that moves by more than 0.1 Å per event. This suggests that each event is localized, typically involving a cutoff radius of less than 7 Å. The microstructural changes primarily occur within a few neighboring cells, indicating a localized relaxation process. The relationship between the aforementioned parameters and the number of randomly displaced atoms at the start of ARTn searches (which depends on the cutoff and the central atom) was not investigated in this works.
@@ IV Conclusions
In this paper, we investigated a-Si using ARTn coupled with a MTP. Seven models of varying sizes were constructed. They displayed excellent agreement with both experimental data and data obtained using a state-of-the-art 100,000-atom MD-based model leveraging a different MLIP. The proposed ARTn-MTP optimizer can further refine and relax configurations that have already been previously optimized using ARTn with a semi-empirical model. Additionally, the relaxation level of a-Si generated through standard molecular dynamics simulations via the melt-quench process can also be improved using the ARTn-MTP coupling scheme. We have generated high-quality amorphous structures of various sizes, including 216, 512, and 1000 atoms, without any traces of local crystallinity. We also found that starting the ARTn simulation with a randomly filled configuration results in an optimized structure with less crystallinity compared to initiating the simulation with an MD prerelaxed configuration. The question of local crystallinity, its link with defects and the ideal amorphous structure has not been addressed previously, due to the lack of good quality models generated through different pathways (non-dynamic, such as with ARTn, and melt-and-quench, for example). The results presented here raise numerous related research questions that could further enhance our understanding of a-Si and its fabrication. Future research could focus on increasing system sizes while minimizing coordination and ring defects, for example, by targeting over-coordinated, under-coordinated, and high-energy atoms during event searches. Further studies will investigate the origins of crystallinity, explore its formation during the preparation of amorphous silicon, and examine methods for producing large-scale amorphous a-Si free of crystallinity.
It is now clear that ARTn coupled with MLIPs is a very effective approach to model a-Si. However, if the primary goal is to replicate experimental conditions as accurately as possible, further innovation is necessary. Experimentally, a-Si is produced mainly by chemical vapor deposition and ion implantation; representative elementary volumes range from scales of tens of nanometers to micrometers, and processing times on the order of seconds to hours. Accessing such time- and length-scales would enable both applied studies–e.g. device-scale simulations–and fundamental studies–e.g. assessing the degree of hyperuniformity of a-Si. Although modeling large systems is straightforward thanks to parallel computing, reaching long timescales remains an elusive challenge. While there are examples of ARTn-based techniques reaching such timescales in irradiated c-Si and a-Si 88, 89, 90, 91, 92, further development is necessary in order for these methods to routinely achieve such outcome.

### v2

@@ III.5 Potential energy
Fig. 9: a) Excess energy between ART-MTP configurations (a-Si) and crystalline silicon (c-Si) as a function of the number of accepted events. A total of 4 events per atom was required to achieve 100% 4-fold coordination in 216-R2 a-Si model. b) Excess energy for models generated using MD simulation solely
@@
Fig. 9 a shows the relationship between energy decrease and the number of events for the model 216-R2. As can be seen, the energy decreases as the number of events increase. A detailed analysis of the events and key parameters is provided in the Nature of the Relaxation section. The experimental excess energy interval is indicated by the light red shaded region in the figure for reference. The inset of Fig. 9 a shows the excess energy per atom for each of our models. These values are also presented in table 4. All models fall within the experimental excess energy interval 88, as detailed in the inset. Notably, the 216-R2 model, which initially contained over 70% coordination defects, was efficiently relaxed to achieve 100% 4-fold coordination and the lowest excess energy of 0.17 eV/atom with just 4 events per atom. To our knowledge, with the exception of a few Keating-potential-based WWW models 10, 85, 29, no other method, including DFT 93, has yielded a model of a-Si with no coordination defects. Fig. 9 b shows the evolution of excess energy as a function of simulation time, with the inset corresponding to the excess energy of the generated a-Si models. Once again, the excess energy falls within the experimental range, although the values are higher than those for the ARTn-MTP-generated a-Si models. The mean excess energy for the ARTn-MTP models is 0.1667 eV per atom, with a spread of 0.0130 eV per atom, while for the MD-generated models, the mean excess energy is 0.1905 eV per atom, with a spread of 0.0035 eV per atom. This demonstrates that the ARTn-MTP-generated a-Si models are more stable than those generated by MD-MTP.
@@ III.6 Density
We computed the density of our models at 0 K. The final density was obtained by running a 50 ps molecular dynamics simulation, cooling the system from 300 K to 0.5 K, and then holding it at 0.5 K for an additional 50 ps, while barostatting the system at 0 Pa. In separate simulations, we also equilibrated each of the aforementioned models at 300 K and 1 atm for 100 ps. The results are presented in table 4. Recall that the calculated density and experimental density of crystalline Si are 2.28 g/cm³ (variable cell relaxation 0 K) and 2.33 g/cm³ (300 K), respectively. The calculated density was obtained using an MTP potential that was trained on DFT data ( 0 K calculations). Therefore, a reasonable comparison between the two densities could not be made due to the difference in temperature conditions. We recorded a mean deviation of 1.46% from the calculated value and 3.57% from the experimental value of crystalline silicon. The MTP potential is trained using a generalized gradient approximation exchange-correlation functional, which is known to overestimate the density of crystalline silicon by 2 % 94. Retraining the MTP on a higher-accuracy exchange-correlation functional, such as the restored regularized strongly constrained and appropriately normed meta-generalized gradient approximation 95 is likely to solve this small discrepancy with experimental values.
@@ III.7 Nature of the relaxation
The primary goal is not only to generate high-quality amorphous structures but also to understand the underlying physics of the relaxation process in amorphous systems over time. Thus, to understand the relaxation process in our amorphous system, we analyze simulation data, focusing on key parameters such as barrier distribution, atomic displacements, active atoms, and energy decrease, considering only accepted events as detailed in Fig. 10. These parameters can help provide insights into the nature and mechanism of the relaxation process. The data were collected under normal deformation, excluding shear deformation during the simulation. We first examine the correlation between atomic displacement and energy barriers. A clear correlation is observed: larger displacements are typically associated with higher energy barriers and more active atoms, as shown in Fig. 10 a. To further validate this, we calculate the correlation between the square of the atomic displacement and the energy barriers. According to Hooke’s law, we expect a quadratic relationship between energy and displacement. When we plot energy against the square of the displacement, we expect a linear relationship, which our results confirm, yielding a strong correlation. A fit was performed to identify potential trends linking atomic displacement to the system’s energy. For instance, Hooke’s law is often applied in similar contexts, where atomic displacement within a certain range is used to determine the spring constant. In this case, a log-log fit indicates a power law exponent of 2.4 with a standard error of ±0.7, considering only the ARTn events in the 4,096-atom system. An analysis over all the ARTn events generated in this study reveals a power law exponent of 1.90 with a standard error of ±0.04. These results suggest that Hooke’s Law likely applies. Next, we find no significant correlation between the energy barriers and the energy gain, nor between energy gain and the number of active atoms, which is consistent with previous studies 96. Furthermore, no correlation exists between the number of atoms displaced up to the saddle point and the final state. As illustrated in Fig. 10 d, a correlation exists between energy barriers and the number of active atoms, with higher energy barriers typically involving a greater number of active atoms. On average, the number of active atoms varies from 10 to 30, with a mean of around 20 atoms that moves by more than 0.1 Å per event. This suggests that each event is localized, typically involving a cutoff radius of less than 7 Å. The microstructural changes primarily occur within a few neighboring cells, indicating a localized relaxation process. The relationship between the aforementioned parameters and the number of randomly displaced atoms at the start of ARTn searches (which depends on the cutoff and the central atom) was not investigated in this works.
@@ IV Conclusions
This paper investigates a-Si using ARTn coupled with a MTP. Seven models of different sizes were developed, showing excellent agreement with both experimental data and a 100,000-atom MD-based model using a different MLIP. The ARTn-MTP optimizer refines configurations optimized by ARTn with a semi-empirical model. It also improves the relaxation of a-Si generated through the melt-quench process. High-quality amorphous structures with 216, 512, and 1000 atoms were produced, all free of local crystallinity. Starting ARTn simulations with a randomly filled configuration yields less crystallinity compared to using an MD-prerelaxed configuration. Overall, ARTn-MTP outperforms MD-MTP in modeling a-Si, as MD-generated a-Si often shows significant crystallinity, whereas ARTn-MTP produces a-Si with minimal crystallinity and low coordination defects.
The relationship between local crystallinity, defects, and ideal amorphous structures remains underexplored due to a lack of high-quality models from different pathways like ARTn and melt-quench. Our results highlight new research directions for improving a-Si understanding and fabrication. Future work could focus on increasing system sizes while minimizing defects by targeting over- and under-coordinated atoms. Additional studies will investigate crystallinity’s origins, its formation in a-Si, and methods for producing large-scale amorphous a-Si without crystallinity.
While ARTn combined with MLIPs is effective for modeling a-Si, further innovation is needed to replicate experimental conditions more accurately. a-Si is mainly produced by chemical vapor deposition and ion implantation, with representative volumes ranging from tens of nanometers to micrometers and processing times from seconds to hours. Achieving these time- and length-scales is essential for both applied and fundamental studies. Despite advancements in parallel computing, reaching long timescales remains a challenge.
Data availability
The potential file and the a-Si models are publicly available at https://gitlab.com/Kazongogit/MTPu/-/tree/main/ART/Si?ref_type=heads. The MTP source code can be found at https://mlip.skoltech.ru/, and the ARTn source code is freely available through a request to Normand Mousseau.